A researcher finds that the mean value of depression = 15, and the variance = 9. Which Depression Score corresponds to a z-score of +1?
Answer #1

Given :-

Mean (\mu) = 15

Variance (\sigma^2) = 9

standard deviation. (\sigma) = \sqrt(\sigma^2) = \sqrt(9) = 3

Let, z-score is z=1

Therefore,

z = \frac {x-\mu}{\sigma}

=> x = \mu + (z*\sigma)

=> x = 15 + (1*3)

x = 18

Answer :- Depression score is 18.